40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3659     Entry time: Wed Oct 6 12:00:23 2010
Author: josephb, yuta, kiwamu 
Type: Update 
Category: CDS 
Subject: Found and fixed filter sampling rate problem with suspensions 

While we looking using dtt and going over the basics of its operation, we discovered that the filter sample rates for the suspensions were still set to 2048 Hz, rather than 16384 Hz which is the new front end.  This caused the filters loaded into the front ends to not behave as expected.

After correcting the sample rate, the transfer functions obtained from dtt are now looking like the bode plots from foton.

We fixed the C1SUS.txt and C1MCS.txt files in the /opt/rtcds/caltech/c1/chans/ directory, by changing the SAMPLING lines to have 16384 rather than 2048.

ELOG V3.1.3-